# MA-LoT: Multi-Agent Lean-based Long Chain-of-Thought Reasoning enhances Formal Theorem Proving

### v2

@@ 2.2.2 NL Long CoT Training
In the second stage, we train a normal instruction-finetuned model to acquire Long CoT reasoning for general NL tasks. We use the OpenO1-SFT-Pro dataset provided by  Open-Source-O1 (2024), a 126k records dataset for general NL question-answering on math and science topics with Long CoT for training. We apply standard next-token prediction SFT, guiding the model to produce Long CoT before it outputs final answers. Throughout the NL Long CoT training, we set the system prompt as follows to explicitly instruct the model to use the Long CoT approach:
You are a helpful assistant who will solve every problem **WITH** Long Chain-of-Thought This prompt effectively “switches on” Long CoT reasoning. The training input includes the system prompt and NL question, with the expected output being the Long CoT and final answer. After training, we observe that the model gains robust NL Long CoT capabilities, which serve as a base for agents to analyze and interact with Lean. However, when applied to Lean4 reasoning, it tends to provide only NL solutions rather than outputting Lean4 code in its output section, indicating the need for further alignment.
@@ 2.2.3 Field-specific alignment
In the final stage of the training process of LoT-TL pipeline, we train the model to make Lean4 Long CoT ability emerge. This is achieved in the training stage by switching to a different system prompt that indicates the model not to use Long CoT and to use a placeholder for the Long CoT content to keep the Long CoT backbone, while not requiring providing any actual Long CoT data in Lean4. Specifically, the system prompt is:
@@
The user asks not to solve with Long CoT, so I will directly write the answer. Using this setup, we first train on the LoT-ProveData for formal theorem proving ability, then train on the LoT-CorrectionData to make the model learn error-analysis and correction skills. The example of training data can be found in Appendix E. We also adopt the curriculum learning data sorting method from Wang et al. (2024) to stabilize training. After training, we find that the Long CoT Lean4 proving and error analysis ability emerges in the LLMs when using the system prompt to turn on Long CoT in inference. We conclude the effectiveness of the TL framework because it preserves the structure of Long CoT and enables the model to activate such capability when instructed.
Following these training steps, we obtain LoT-Solver, which has the emergent Lean reasoning capability in Long CoT from data without Lean Long CoT annotation. This makes it able to perform the roles of both whole-proof prover and tree-search corrector with deep NL-FL joint thinking.
2.3 Multi-Agent FL Proof Writing
@@
This section presents (to the best of our knowledge) the first multi-agent framework that combines the advantages of whole-proof generation and tree-search methods under the Lean-based Long CoT paradigm. We use the LoT-Solver as the base model for both agents. Under this setup, we use the prover agent to write a complete proof draft (Section 2.3.1) and apply the corrector agent to analyze and correct the proof based on Lean verifier feedback (Section 2.3.2).
2.3.1 Prover Agent
@@
The Prover agent writes the initial Lean4 proof using a whole-proof generation strategy, then submits it to the Lean4 verifier to check the correctness and passes it to the corrector agent if the proof is wrong. We use the system prompt to turn on the Long CoT reasoning and use a specific header in Long CoT to guide the model to make a high-level proof plan. Here is the instruction template for the prover agent:
@@
{... **WITH** Long CoT ...}### Instruction:{NL statement}{FL statement}### Response:Alright, I should do the following:1. Provide the natural language analysis for the theorem based on the Natural language theorem statement.2. Draft the Lean4 tactics I should use to solve the problem3. Write the output Lean4 code. The complete example of prover agent input and output can be found in Appendix F. The emergent Lean reasoning ability in the Long CoT enables the model can write a better-structured proof based on its high-level plan compared to direct proof generation. Upon generating the proof, the prover agent submits it to the Lean evaluator for verification. If incorrect, the theorem is passed to the corrector agent for further refinement.
2.3.2 Corrector Agent
@@
The corrector agent functions like the tree-search method. Upon receiving a wrong proof and Lean verifier feedback, it systematically analyzes them in Long CoT. Then, after re-evaluating the proof strategy and rethinking, the model generates a revised proof that intends to solve the errors.
@@
The instruction prompt remains identical to the prover agent. We incorporate the incorrect proof and feedback from the Lean verifier in the Long CoT followed by instructions to direct the model to analyze the error and formulate a revised proof. Detailed examples of such prompts are available in Appendix F. Then, the corrector agent will pass the new proof to the Lean4 verifier, if the new proof is still wrong, iteratively analyze the errors until success or reach the max retry limit.
@@
The corrector agent enhances theorem proving by enabling deeper reflection and systematic exploration of alternative proof strategies based on error messages. This iterative correction process increases the likelihood of discovering non-trivial proofs while maintaining computational efficiency.
@@ 3 Experiments
We conduct comprehensive experiments on the MiniF2F-Lean4 (Zheng et al., 2021) dataset to assess the performance of MA-LoT framework in formal proof writing. Specifically, we evaluate the superiority of our multi-agent Long CoT system in writing better-structured and more insightful proofs by additional proof it could fine under similar sampling budget (Section 3.3) Additionally, we perform studies on our corrector agent (Section 3.4), training components of LoT-Solver (Section 3.5), and case study (Section 3.6) to further analyze the impact of individual components.
@@
Table 1: Main experimental results of MA-LoT. Our results is presented as Base model + method. The LoT (whole-proof) indicates using the whole-proof results of our LoT-Solver model and MA-LoT indicates the result of our whole pipeline. The sampling budget $x+k\times y$ in the MA-LoT framework indicates we first perform $x$ whole-proof writing using the prover agent and $k$ rounds of correction using the corrector agent. In practice, one round of correction costs approximately half of the whole-proof generation budget, result in $y=\frac{1}{2}x$.
@@ 3.1.1 Dataset and Task
Both the test and validation datasets contain 244 Lean4 statements. The range of problems varies from high-school competition questions to undergraduate-level theorem proofs. It contains 488 problems from three sources: (1) 260 problems sampled from the MATH dataset (Hendrycks et al., 2021); (2) 160 problems from high-school math competitions including AMC, AIME, and AMO; (3) 68 manually crafted problems at the same difficulty level as (2). Our task is to query the LLM to generate Lean4 proofs for MiniF2F problems based on their formal statements and NL descriptions. To minimize computational overhead, imports are manually configured.
@@ 3.1.2 Baselines
To highlight MA-LoT ’s capabilities, we select some of the most competitive baselines in recent years, covering both tree-search and whole-proof generation approaches. For tree-search methods, we include: Expert Iteration (Polu et al., 2022), Llemma (Azerbayev et al., 2023), ReProver (Yang et al., 2024b), Lean-STaR(Lin et al., 2024), and InternLM2.5-StepProver(Wu et al., 2024a) as our baselines. For whole-proof generation baselines, we include closed-source LLMs, represented by GPT-4-Turbo(Achiam et al., 2023) and Gemini-1.5(Reid et al., 2024), and open-source expert models, such as DeepSeek-Math(Shao et al., 2024), TheoremLlama(Wang et al., 2024), DeepSeek-Prover-v1.5-RL (Xin et al., 2024b), STP-Lean (Dong & Ma, 2025), and Godel-Prover-SFT (Lin et al., 2025).
For whole-proof generation baselines, we set the sample budget to pass@128 with 4096 context length (expect for Godel-Prover, where, following the original paper, we use pass@32), balancing robustness and manageable GPU consumption. For tree-search methods, we align the search cost as closely as possible to whole-proof generation. [^2] This explains why we exclude RMaxTS for DeepSeek-Prover-v1.5, as its smallest disclosed sample budget ($1\times 3200$) not a comparable result to our baselines.
@@ 3.2 Implementation Details
In the training process of the model, we use Openo1-SFT-Pro, LoT-ProverData, and LoT-CorrectionData to train two base models, namely DeepSeek-Prover-v1.5-SFT and Godel-Prover-SFT. For different training stages, the learning rate is as follows: 1E-5 for NL Long CoT training, 1E-7 for LoT-TL on LoT-PD, and 1E-6 for LoT-CD. The total computational cost for training is around 1 GPU day, and evaluation is 11 GPU days on a 4×4$\times$4 × H100-96G cluster [^3] The high evaluation cost due to accelerate inference methods like vLLM unable to fit our machine. To evaluate our framework in detail, we present three sets of results, including: (1) LoT (whole-proof): pass@128 whole-proof generation result of LoT-Solver. (2) MA-LoT: Our primary evaluation result, where the prover performs 64 whole-proof generations and undergoes two rounds of corrector refinement [^4] Because we don’t need to pass correct proof to corrector, two rounds of correction is approximately the same as one round of whole-proof generation. (3) Cumulative Results: A combined evaluation aggregating all LoT models outputs obtained throughout the experiment process.
Our code, model, and data will be published in https://github.com/RickySkywalker/LeanOfThought-Official
3.3 Results
Table 1 presents our main results, showing that MA-LoT achieves 61.07% accuracy rate on MiniF2F-Test benchmark and 57.79% for LoT-Solver using whole-proof generation. After the enhancement of MA-LoT, we achieve 10.37% on Godel-Prover and 12.72% on DeepSeek-Prover on the MiniF2F-Test benchmark. This significant and uniform improvement of our framework indicates its effectiveness. Detailed analysis also shows that our models can solve IMO and AIME problems that previous models struggled with. Our model surpasses state-of-the-art tree-search (InternLM-2.5) and whole-proof generation (Godel-Prover) baselines, demonstrating that our proposed multi-agent framework based on Long CoT excels in formal theorem proving.
MA-LoT outperforms all tree-search baselines by at least 20.45% because its prover agent constructs proosfs with high-level NL planning using emergent Lean Long CoT reasoning capability. This indicates our prover agent can leverage LLMs’ strong NL reasoning ability, which leads to more comprehensive proofs. Additionally, MA-LoT surpasses whole-proof generation baselines by at least 10.37%, as its corrector agent analyzes, reflects, and reformulates proofs based on Lean4 executor feedback in Long CoT. The strong performance also demonstrates the effectiveness of our ideas of integrating FL verification in NL Long CoT reasoning with its emergent capability. Notably, the DeepSeek-Prover + MA-LoT is based on DeepSeek-Prover-v1.5-SFT; it outperforms its RL-trained variant by 6.15%. This suggests that our multi-agent training framework and Lean-based Long CoT methodology align more naturally with formal theorem proving than RL alone.
Method Prover 1 round 2 round 3 round
@@ 3.3 Results
Godel 54.92% 59.43% 61.07% 61.89%
Table 2: Result of different rounds of corrector agent correction with DeepSeek-Prover and Godel-Prover as base models
@@
Table 3: Ablation study result in pass@64 under DeepSeek-Prover as base model
@@
The comparison between LoT (whole-proof) and MA-LoT further highlights the importance of our multi-agent framework. We observe a 2.46% improvement by reallocating the computation resources from the prover agent doing more whole-proof generation to the corrector agent to analyze and refine proof based on Lean executor feedback. This validates the necessity of an iterative prover-corrector multi-agent system over relying solely on a prover agent. In summary, these results confirm that Long CoT reasoning, combined with formal verification and a multi-agent paradigm, enhances the discovery of non-trivial and in-depth proofs, thereby validating the effectiveness of our proposed method.
3.4 Corrector Agent study
@@
To assess the impact of the correct agent in MA-LoT, we present the cumulative accuracy on the MiniF2F-Test across different rounds of correction in Table 2 for two of our base models. The prover column presents the pass@64 (or 16) accuracy rate of the prover agent in whole-proof generation, while Round-$i$ columns indicate successive correction rounds. The results indicate that during the first three correction rounds, the corrector agent successfully correct an average of 11.55% of theorems that the prover unable to answer. Our analysis shows that most corrected proofs belong to IMO, AIME, and high-level MATH problems, which are particularly difficult for prior models. This highlights the corrector agent’s ability to analyze feedback from the Lean4 executor feedback using emergent Lean capability in Long CoT to discover non-trivial proofs. The case study for the analysis and regeneration of new proofs can be found in Section 3.6 and Appendix D.

### v3

@@ 2.2.2 NL Long CoT Training
In the second stage, we train a normal instruction-finetuned model to obtain the general Long CoT reasoning capability in Natural Language. We use the OpenO1-SFT-Pro dataset provided by  Open-Source-O1 (2024), a 126k records dataset for general NL question-answering on math and science topics with Long CoT for training. We apply standard next-token prediction SFT, guiding the model to produce Long CoT before it outputs final answers. Throughout the NL Long CoT training, we set the system prompt as follows to explicitly instruct the model to use the Long CoT approach:
You are a helpful assistant who will solve every problem **WITH** Long Chain-of-Thought This system prompt indicates that the model should use the Long CoT to write the answer. The training input includes the system prompt and NL question, with the expected output being the Long CoT and final answer. After training, we observe that the model gains robust NL Long CoT capabilities, which serve as a base for models to analyze and interact with Lean code. However, when applied to Lean4 reasoning, it tends to provide only NL solutions rather than outputting Lean4 code in its output section, indicating the need for further alignment.
@@ 2.2.3 Field-specific alignment
In the final stage of the training process of LoT-TL pipeline, we train the model to obtain emergent Lean4 Long CoT ability. In this training stage, we set the system prompt to instruct the model not to use Long CoT when answering. In the meantime, to make the model aware of the Long CoT structure, we use a simple placeholder to occupy the original place of the Long CoT. The above method allows us to train the model aware of the Lean4 Long CoT structure without requiring any Lean4 Long CoT data. Specifically, the system prompt is:
@@
The user asks not to solve with Long CoT, so I will directly write the answer. Under this setup, we first train on the LoT-ProveData for formal theorem proving ability, then train on the LoT-CorrectionData to make the model learn error-analysis and correction skills. The example of training data can be found in Appendix F. We also adopt the curriculum learning data sorting method from Wang et al. (2024) to stabilize training. After training, we find that the Long CoT Lean4 proving and error analysis ability emerges in the LLMs when using the system prompt to turn on Long CoT in inference. We conclude the effectiveness of the TL framework because it preserves the structure of Long CoT and enables the model to activate such capability when instructed.
Following the above steps, we train the LoT-Solver model based on an expert Lean4 prover with instruction fine-tuning. LoT-Solver is a model with Lean4 Long CoT reasoning capability, which also enables the separation of whole-proof writer and corrector.
2.3 Model-Collaboration Lean4 Proof Writing
@@
This section presents the model-collaboration framework that combines the advantages of whole-proof generation and tree-search methods under the Lean-based Long CoT paradigm. We use the LoT-Solver as the base model for both the prover and the corrector. Under this setup, we use the prover model to write a complete proof draft (Section 2.3.1) and apply the corrector model to analyze and correct the proof based on Lean verifier feedback (Section 2.3.2).
2.3.1 Prover Model
@@
The prover model writes the initial Lean4 proof using a whole-proof generation strategy. Then, it is submitted to the Lean4 verifier to check its correctness and passed to the corrector model if it is wrong. We use the system prompt to turn on the Long CoT reasoning and use a specific header in Long CoT to guide the model in making a high-level proof plan. Here is the instruction template for the prover model:
@@
{... **WITH** Long CoT ...}### Instruction:{NL statement}{FL statement}### Response:Alright, I should do the following:1. Provide the natural language analysis for the theorem based on the Natural language theorem statement.2. Draft the Lean4 tactics I should use to solve the problem3. Write the output Lean4 code. The prover model’s input and output example can be found in Appendix G. The emergent Lean reasoning ability in the Long CoT enables the model to write a better-structured proof based on its high-level plan than direct proof generation. Upon generating the proof, the prover model submits it to the Lean evaluator for verification. The theorem is then passed to the corrector model for further refinement if incorrect.
2.3.2 Corrector Model
@@
After receiving a wrong proof and Lean verifier feedback, the corrector model will systematically analyze them in Long CoT. The corrector re-evaluates and rethinks the proof strategy, then generates a revised proof that intends to correct the errors and complete the proof. In conceptual analogy, the corrector model functions similarly to the tree-search method but with greater flexibility and deeper analysis.
@@
The instruction prompt remains identical to the prover model. We incorporate the incorrect proof and feedback from the Lean verifier in the Long CoT, followed by instructions to direct the model to analyze the error and formulate a revised proof. Detailed examples of such prompts are available in Appendix G. Then, the corrector model will pass the new proof to the Lean4 verifier. If the new proof is still wrong, iteratively analyze the errors until success or reach the max retry limit.
@@
The corrector model enhances theorem proving by enabling deeper reflection and systematically exploring alternative proof strategies based on error messages. This iterative correction process increases the likelihood of discovering non-trivial proofs while maintaining computational efficiency.
@@ 3 Experiments
We conduct comprehensive experiments on the MiniF2F-Lean4 (Zheng et al., 2021) dataset to assess the performance of the MA-LoT framework in formal proof writing. Specifically, we intend to prove the advantage of the MA-LoT framework by showing its capability to generate better-structured and more insightful proofs. We do this by showing our model has a better general performance in Section 3.3. Moreover, we perform studies on our corrector model in Section 3.4, the efficiency of the Long CoT method in Section 3.5, training components of LoT-Solver in Section 3.6, and a case study in Section 3.7 to further analyze the impact of individual components.
@@
Table 1: Main experimental results of MA-LoT. Our results are presented as Base model + method. The LoT (whole-proof) indicates using the whole-proof results of our LoT-Solver model, and MA-LoT indicates the result of our whole pipeline. The sampling budget $x+k\times y$ in the MA-LoT framework indicates we first perform $x$ whole-proof writing using the prover model and $k$ rounds of correction using the corrector model. In practice, one round of correction costs approximately half of the whole-proof generation budget, resulting in $y=\frac{1}{2}x$.
@@ 3.1.1 Dataset and Task
Both the test and validation datasets contain 244 Lean4 statements. The range of problems varies from high-school competition questions to undergraduate-level theorem proofs. It includes 488 problems from three sources: (1) 260 problems sampled from the MATH dataset (Hendrycks et al., 2021); (2) 160 problems from high-school math competitions, including AMC, AIME, and AMO; (3) 68 manually crafted problems at the same difficulty level as (2). Our task is to query the LLM to generate Lean4 proofs for MiniF2F problems based on their formal statements and NL descriptions. To minimize computational overhead, imports are manually configured.
@@ 3.1.2 Baselines
To highlight MA-LoT ’s capabilities, we select some of the most competitive baselines in recent years, covering both tree-search and whole-proof generation approaches. For tree-search methods, we include: Expert Iteration (Polu et al., 2022), Llemma (Azerbayev et al., 2023), ReProver (Yang et al., 2024b), Lean-STaR(Lin et al., 2024), and InternLM2.5-StepProver(Wu et al., 2024a) as our baselines. For whole-proof generation baselines, we include closed-source LLMs, represented by GPT-4-Turbo(Achiam et al., 2023), Gemini-1.5(Reid et al., 2024), and DeepSeek-V3 (Liu et al., 2024). Moreover, we include representative open-source expert models for whole-proof writing, such as DeepSeek-Math(Shao et al., 2024), TheoremLlama(Wang et al., 2024), DeepSeek-Prover-v1.5-RL (Xin et al., 2024b) [^2] The reported results of DeepSeek-Prover-V1.5 and Goedel-Prover in our paper are different from the original paper because we are unable to access vLLM in our machine, and it is known that the inference without vllm will lead to slight drop of the model’s capability, STP-Lean (Dong & Ma, 2025), and Godel-Prover-SFT (Lin et al., 2025).
For whole-proof generation baselines, we set the sample budget to pass@128 with 4096 context length (except for Godel-Prover, where, following the original paper, we use pass@32), balancing robustness and manageable GPU consumption. We align the search cost as closely as possible for tree-search methods to whole-proof generation. [^3] This explains why we exclude RMaxTS for DeepSeek-Prover-v1.5, as its smallest disclosed sample budget ($1\times 3200$) is not a comparable result to our baselines.
@@ 3.2 Implementation Details
Method Prover round 1 round 2 round 3
@@
Godel-Prover-SFT 54.92% 59.43% 61.07% 61.89%
Prover as Corrector 54.92% 56.15% 57.38% -
Table 2: Results from multiple rounds of correction using DeepSeek-Prover and Godel-Prover as base models, along with the outcomes of using the prover to perform correction.
@@
Table 3: Ablation study result in pass@64 under DeepSeek-Prover as base model.
Method Godel-Prover-SFT MA-LoT-Godel
Sample budget pass@32 $16+2\times 8$
Avg. tokens gen 492.10 657.54
MiniF2F-Test 58.20% 61.07%
Table 4: Long CoT efficiency study result.
In the model’s training process, we use Openo1-SFT-Pro, LoT-ProverData, and LoT-CorrectionData to train two base models, namely DeepSeek-Prover-v1.5-SFT and Godel-Prover-SFT. For different training stages, the learning rate is as follows: 1E-5 for NL Long CoT training, 1E-7 for LoT-TL on LoT-PD, and 1E-6 for LoT-CD. The total computational cost for training is around 1 GPU day, and evaluation is 11 GPU days on a 4×4$\times$4 × H100-96G cluster [^4] The high evaluation cost is due to accelerated inference methods like vLLM being unable to fit our machine. To evaluate our framework in detail, we present three sets of results, including: (1) LoT (whole-proof): pass@128 whole-proof generation result of LoT-Solver. (2) MA-LoT: Our primary evaluation result, where the prover performs 64 whole-proof generations and undergoes two rounds of corrector refinement [^5] Because we don’t need to pass a correct proof to the corrector, two rounds of correction are approximately the same as one round of whole-proof generation. (3) Cumulative Results: A combined evaluation aggregating all LoT models’ outputs obtained throughout the experiment process.
3.3 Results
Table 1 presents our main results, showing that MA-LoT achieves 61.07% accuracy rate on MiniF2F-Test benchmark and 57.79% for LoT-Solver using whole-proof generation. After the enhancement of MA-LoT, we achieve 10.37% on Godel-Prover and 12.72% on DeepSeek-Prover on the MiniF2F-Test benchmark. This significant and uniform improvement of our framework indicates its effectiveness. Detailed analysis also shows that our models can solve IMO and AIME problems that previous models struggled with. Our model surpasses state-of-the-art tree-search (InternLM-2.5) and whole-proof generation (Godel-Prover) baselines, demonstrating that our proposed model-collaboration framework based on Long CoT excels in formal theorem proving.
MA-LoT outperforms all tree-search baselines by at least 20.45% because its prover model constructs proofs with high-level NL planning using emergent Lean Long CoT reasoning capability. This indicates our prover model can leverage LLMs’ strong NL reasoning ability in its Long CoT, which leads to more comprehensive proofs. Additionally, MA-LoT surpasses whole-proof generation baselines by at least 10.37%, as its corrector model analyzes, reflects, and reformulates proofs based on Lean4 executor feedback in Long CoT. The strong performance also demonstrates the effectiveness of our ideas of integrating FL verification in NL Long CoT reasoning with its emergent capability. Notably, the DeepSeek-Prover + MA-LoT is based on DeepSeek-Prover-v1.5-SFT; it outperforms its RL-trained variant by 6.15%. This suggests that our model-collaboration framework and Lean-based Long CoT methodology align more naturally with formal theorem proving than RL alone.
@@ 3.3 Results
The comparison between LoT (whole-proof) and MA-LoT further highlights the importance of our model-collaboration framework. We observe a 2.46% improvement by reallocating the computation resources from the prover model, which performs more whole-proof generation, to the corrector model, which analyzes and refines proofs based on Lean executor feedback. This validates the necessity of a model-collaboration system over relying solely on a prover model. In summary, these results confirm that Long CoT reasoning, combined with formal verification and a model-collaboration paradigm, enhances the discovery of non-trivial and in-depth proofs, thereby validating the effectiveness of our proposed method.
3.4 Corrector study
@@
To assess the impact of the correct model in MA-LoT, we present the cumulative accuracy on the MiniF2F-Test across different rounds of correction in Table 2 for two of our base models. The prover column presents the pass@64 (or 16) accuracy rate of the prover model in whole-proof generation, while Round-$i$ columns indicate successive correction rounds. The results indicate that during the first three correction rounds, the corrector model successfully corrects an average of 11.55% theorems that the prover cannot answer. Our analysis shows that most corrected proofs belong to IMO, AIME, and high-level MATH problems, which are particularly difficult for prior models. Such additional solved problems highlight the corrector’s ability to analyze feedback from the Lean4 executor feedback using emergent Lean capability in Long CoT to discover non-trivial proofs. The case study for the analysis and regeneration of new proofs can be found in Section 3.7 and Appendix E.
Additionally, to validate the effectiveness of the expert corrector, we conducted an experiment that used prover to perform corrections. In particular, we apply the LoT-Solver based on Goedel-Prover-SFT. We add the wrong Lean4 proof and error message as a comment in the Lean4 code block and ask the prover to correct it. Results are shown in the last row of Table 2. It shows a clear drop in performance in both correction rounds. This drop is because the prover model tends to directly repeat the wrong proof instead of using the Long CoT to analyze the error and correct the proof. This experiment shows the importance of an expert corrector model, further validating the effectiveness of our idea of using FL as a backbone for Long CoT in NL.
3.5 Long CoT Efficiency Study
To further study the property of Long CoT, we perform this experiment to test whether MA-LoT is able to search for the answer more efficiently. In this experiment, we measure the GPU hours as well as the tokens generated in both MA-LoT and the base model. We observed that the MA-LoT generated 33% additional tokens and costs 70% additional GPU hours compared to the base model. Thus, to make a fair comparison in the sense of computation hours, we enlarge the base model’s sample budget to 1.7 times in this experiment; the results are presented in the Tab. 4. The table shows that despite having a smaller sampling budget, MA-LoT still outperforms the base model by 4.93%, which indicates that MA-LoT is a better way to allocate computation power.